What do you understand by bond energy term?

ii) Use the following to calculate the Cβˆ’H and πΆβˆ’πΆ bond energies.

C2H6(g) yields 2C(g) + 6H(g) Ξ”π»βˆΆ282 πΎπ½π‘šπ‘œπ‘™βˆ’1

C3H6(g) yields 3C(g) + 8H(g) Ξ”π»βˆΆ3996 πΎπ½π‘šπ‘œπ‘™βˆ’1


1
Expert's answer
2021-02-01T03:41:50-0500

Bond energy is the amount of energy required to break a chemical bond. It is also the amount of energy released when a chemical bond is formed.


From equation 1, we have 6 C-H, 1 C-C and 3 H-H bonds. Let C-H bond be x, C-C bond be y and H-H bond be z.

βˆ†H= Hp - Hr

282= [y + 3z] -6x.......1

From eq2, we have 8C-H bonds, 1.5 C-C Bonds and 4H-H bonds.

3996= 1.5y + 4z - 8x.......2

Since we are only interested in C-H and C-C bonds, let's eliminate H-H. Multiply (1) by 4 and (2) by 3

1128= 4y + 12z -24x.....(3)

1198= 4.5y + 12z -24x

Subtract

-10860 = -0.5y

y= 21720

In the Elimination, z=0, substitute 0 for z and 21720 for y in (1)

282= 21720 + 3(0) -6x

282-21720 = -6x

-21438= -6x

x= -21438/6

x= 3573


Therefore

x= C-H bond energy = 3573Kj/mol

y= C-C bond energy = 21720Kj/mol